OpenStudy (anonymous):

PLEASEE HELP! MEDALS GIVIN Two trains leave the station at the same time, one heading west and the other east. The westbound train travels at 85 miles per hour. The eastbound train travels at 65 miles per hour. How long will it take for the two trains to be 240 miles apart?

OpenStudy (perl):

we can model this with algebraic equations

OpenStudy (perl):

eastbound train: x1(t) = 65*t westbound train : x2(t) = -85*t distance between them is x1(t) - x2(t) = 65*t - (-85*t) = 150*t

OpenStudy (anonymous):

Two cars leave towns 450 kilometers apart at the same time and travel toward each other. One car's rate is 16 kilometers per hour less than the other's. If they meet in 3 hours, what is the rate of the slower car?